MIDC News
Year of 2004
December 14, 2004
- 14:51 MST: An instrument's cleaning maintenance history since November 4, 2004 is now available for SRRL BMS and BSRN. The records for an instrument can be found by clicking the Maintenace (Maint) link below an instrument model number on the SRRL BMS & BSRN Instrument History Page. -Afshin Andreas
- 14:57 MST: A "DISABLE automatic browser refresh" feature has been added to the MIDC live data pages, just below "Generate Custom Data". By clicking on it you can disable the sometimes annoying browser refresh (to re-enable just click "ENABLE automatic browser refresh").
-Afshin Andreas
- 11:09 MST: November 2004 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 11:01 MST: November 2004 Monthly Data Sets & Reports for SRRL BMS have been generated. -Afshin Andreas
- 08:58 MST: The SRRL TSI-880 will be down for an extended period of time due to the Astrodyne power supply failing again due to a power dip at SRRL yesterday. -Afshin Andreas
- 14:31 MST: The (3) SRRL BMS Wind Speed sensors were replaced today between approx 12:00 to 14:00 MST today. The old 42' sensor had a broken cup and the old 22' sensor had a cracked cup. -Afshin Andreas
- 08:53 MST: The SRRL TSI-880 is currently down to an issue with the charging system/battery, cloud cover is unavailable at this time. -Afshin Andreas
- 13:28 MST: The SRRL BMS "Intrument QA" page has been replaced with a database available through AIM. Known instrument problems and data quality statements (DQS) can be found by clicking the DQS links below the instrument model on the SRRL BMS Instrument History Page. Note: A future link to an instruments cleaning history will be available sometime later. -Afshin Andreas
- 07:45 MST: The correct cal factor for the replacement Global TUVR was updated in the SRRL BMS logger today @ 7:43 MST. -Afshin Andreas
- 16:02 MST: The SRRL BMS Global TUVR S/N 33444 was returned to service on Nov. 18 @ 11:14 MST. The instrument seems to be working now (there was no change/service done to it). -Afshin Andreas
- 07:54 MST: An SPA web-based solar position calculator is now available on the MIDC. The Solar Position Algorithm (SPA) calculates the location of the sun in the sky (and various other values) in the period from the year -2000 to 6000 CE, with uncertainties of +/- 0.0003 degrees. It is more accurate than SOLPOS. -Afshin Andreas
- 15:46 MST: October 2004 Monthly Data Sets & Reports for SRRL BMS have been generated. -Afshin Andreas
- 13:53 MST: The SRRL BMS spare 8-48 was calibrated in BORCAL 2004-01. The instrument was swapped out today and the logger program was updated @ 13:24 MST. -Afshin Andreas
- 10:33 MST: October 2004 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 09:51 MST: The NWTC M2 Dew Point Sensor was replaced on Nov. 5 @ 12:26 MST. Since October 1 (approximate), the old sensor was reading low by about 2-4 degrees resulting in RH values that were about 15-20% low. This problem only seemed to occur during high RH conditions (greater than approx 50%). It is not certain at this time whether the new sensor has fixed the problem. -Afshin Andreas
- 14:13 MST: The SRRL BMS IR instruments (Downwelling PIR & CG4 and Upwelling PIR) were returned to service today between 13:23 and 13:52 MST, after being calibrated in the NREL Black Body Cal System. Cal coeficients were updated in the logger at 13:56 MST. -Afshin Andreas
- 15:59 MST: There was a power failure at SRRL today from 8:18 to 14:42 MST. -Afshin Andreas
- 07:07 MST: The SRRL BMS Downwelling CG4 was swapped with a reference CG4 S/N 10547 yesterday @ 15:46 MST, which was calibrated in the new NREL blackbody system on 10/20/2004. This CG4 will be left in BMS for about a week and then the original CG4 (which will be calibrated in the blackbody) will be put back in BMS. New coefficeints were entered into the logger at 7:00 MST today. -Afshin Andreas
- 14:37 MST: The SRRL Global 40-South Spectroradiometer was removed from service today at 11:50 MST for calibration and it was returned to service today @ 14:25 MST. Prior to calibration the unit was found to be within 2% of NIST irradiance for wavelengths below 1000nm. -Afshin Andreas
- 11:18 MST: The SRRL BMS Upwelling PIR was swapped with a reference PIR S/N 31196F3, which was calibrated in the new NREL blackbody system on 10/21/2004. This PIR will be left in BMS for about a week and then the original PIR (which will be calibrated in the blackbody) will be put back in BMS. New coefficeints were entered into the logger at 11:16 MST today. -Afshin Andreas
- 08:52 MST: The SRRL BMS Downwelling PIR was swapped with a reference PIR S/N 31197F3, which was calibrated in the new NREL blackbody system on 10/20/2004. This PIR will be left in BMS for about a week and then the original PIR (which will be calibrated in the blackbody) will be put back in BMS. New coefficeints were entered into the logger at 8:41 MST today. -Afshin Andreas
- 07:43 MST: ALL SRRL BMS Global/Direct UVA/UVB/UVE Radiometers were calibrated on 9/15/2004, new factors were entered into the logger today @ 7:34 MST. -Afshin Andreas
- 08:36 MST: September 2004 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 13:12 MST: The SRRL BMS Report output for the Zenith, Azimuth, & Airmass was modified for Sept 2004, so that data for all hours is included in the report (instead of just daylight hours). This was done to prevent a bogus discontinuity in the daily averages (in monthly reports & plots) caused by the sunset crossing an hourly boundary. Note: Historical monthly report data may be corrected sometime in the future. -Afshin Andreas
- 13:06 MST: September 2004 Monthly Data Sets & Reports for SRRL BMS have been generated. -Afshin Andreas
- 15:02 MST: The SRRL BMS Global CM3 (CNR1) was calibrated against the Calculated Global from the BMS CH1 and 8-48 on 8-28-2004. The cal factor was updated in the logger today @ 14:45 MST. -Afshin Andreas
- 08:07 MST: The SRRL BMS physical instrument mounting locations on deck and two towers PDF file has been updated to include the 8 vertical instruments installed on the radiometer tower. -Afshin Andreas
- 14:17 MST: Fixed a minor documentation problem with the SRRL BMS report generator that would display "X.X% exceed the 15% QC threshold", when it should have simply said "X.X% failed QC". This only affected instruments that were tested during day light hours (usually irradiance instruments) that were NOT using the SERIQC test. Reports (both PDF and CSV) were only corrected for August 2004. -Afshin Andreas
- 07:32 MST: Calibration Date/Factor was corrected on SRRL BMS Global Zenith Angle Correted PSP Report (both PDF and CSV) from August 2004. -Afshin Andreas
- 13:59 MST: The SRRL BMS Custom Monthly Data for Aug 2004 is now available. -Afshin Andreas
- 10:36 MST: August 2004 Monthly Data Sets & Reports for SRRL BMS have been generated, which include 3 new measurements. Note: Custom monthly data for Aug 2004 is currently unavailable. -Afshin Andreas
- 10:05 MST: August 2004 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 15:21 MST: The following SRRL BMS Radiometer Tower spare pyranometers were calibrated in BORCAL 2004-03: Upwelling PSP, Global 90-East PSP, Global 90-East LI-200, & Global 90-North LI-200. Instruments were swapped out today and the logger program was updated @ 15:01 MST -Afshin Andreas
- 13:41 MST: YESDAQ is operational, no data loss occured. -Afshin Andreas
- 06:48 MST: YESDAQ is currently down due to hard drive crash, therefore SRRL cloud cover is unavailable. -Afshin Andreas
- 15:07 MST: YESDAQ is operational again, the database had to be restored from backup, no data loss occured. -Afshin Andreas
- 13:53 MST: The SRRL TSI-880 data collection has been down since 13:13 MST, due to a corrupt record in the YESDAQ database which cannot be deleted at this time. -Afshin Andreas
- 13:41 MST: The SRRL BMS Global and Direct TUVRs were removed from service today @ 10:51 MST for troubleshooting. -Afshin Andreas
- 13:07 MST: The SRRL BSRN Diffuse PSP which was calibrated in BORCAL 2004-02 was swapped out today, cal factor was updated in the logger @ 13:06 MST. -Afshin Andreas
- 11:11 MST: The following SRRL BMS Radiometer Tower LI-200 spare pyranometers were calibrated in a more recent BORCAL: Global 90-North & Global 90-East. Instruments were swapped out today and the logger program was updated @ 11:08 MST. -Afshin Andreas
- 10:47 MST: The following SRRL BMS Radiometer Tower PSP spare pyranometers were calibrated in BORCAL 2004-02: Global 40-South & Global 90-West. Instruments were swapped out today and the logger program was updated @ 10:28 MST -Afshin Andreas
- 09:40 MST: The following SRRL BMS Radiometer Tower LI-200 spare pyranometers were calibrated in BORCAL 2004-02: Global 40-South, Global 90-South, Global 90-West, Upwelling. Instruments were swapped out today and the logger program was updated @ 9:14 MST. -Afshin Andreas
- 12:37 MST: The SRRL BMS data set since Aug 1 @ 00:00 MST has been updated to include Tower Wet-Bulb Temperature. -Afshin Andreas
- 09:26 MST: Wet Bulb Temperature (Tower) has been added to the SRRL BMS datastream. -Afshin Andreas
- 08:51 MST: The Wet-Bulb Temperature is now available on the SRRL BMS Display. This is a calculation based on the dry-bulb temp, dew point, and local station pressure. The wet-bulb will be added to the actual BMS data stream sometime soon. -Afshin Andreas
- 13:48 MST: The SRRL BMS Global PSP (ventilated) was calibrated in BORCAL 2004-02. The Cal Factor @ 45 and the responsivity coefficents were updated in the data acquisition system today @ 13:44 MST -Afshin Andreas
- 13:16 MST: The following SRRL BMS instruments which were calibrated during BORCAL 2004-02, had their cal factors updated in the logger program today @ 13:09 MST: Globals: CM6b, TSP-700, PSP, TSP-1, SP-Lite, LI-200 #1, LI-200 #2; Directs: CH1, NIP #1, NIP #2, LI-201. -Afshin Andreas
- 09:48 MST: The SRRL BMS data set since Aug 1 @ 00:00 MST has been updated to include the Total & Opaque Cload Cover from the YES TSI-880. -Afshin Andreas
- 07:59 MST: The Opaque Cloud Cover was added to the SRRL BMS data stream today @ 07:48 MST. -Afshin Andreas
- 06:47 MST: The Total Cloud Cover calculated by the SRRL TSI-880 Total Sky Imager was added to the SRRL BMS data stream yesterday @ 15:21 MST. The Opaque Cloud Cover will be added in the near future. -Afshin Andreas
- 08:08 MST: July 2004 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 07:43 MST: July 2004 Monthly Data Sets & Reports for SRRL BMS have been generated. -Afshin Andreas
- 08:13 MST: The Calibration factors for the SRRL BMS Yankee TSP-700 and TSP-1 were updated in the logger today @ 7:49 MST. The TSP-1 is now using the cal factor from BORCAL 2003-03 and the TSP-700 is using a cal factor computed from BMS reference instruments on 7/18/2004. -Afshin Andreas
- 12:22 MST: The SRRL Yankee TSI-880 is now collecting Cloud Cover every 1 minute (instead of 10-minutes). -Afshin Andreas
- 07:31 MST: The SRRL TSI-880 was operational again as of July 23 @ 17:20 MST. -Afshin Andreas
- 12:59 MST: The SRRL Yankee TSI-880 has been shutdown for maintenance. -Afshin Andreas
- 08:11 MST: The SRRL BMS Instantanous Data Display updated every 6 seconds, is now available. Access to the page is restricted to the NREL Mesa Top and OTF. -Afshin Andreas
- 11:12 MST: A Yankee TSI-880 Total Sky Imager was installed at SRRL on July 14, 2004. The Percent Cloud Cover calculated by this instrument is now availabe on the SRRL BMS Display as of 10:09 MST today. Full access to the TSI historical data and images may be available in the future. -Afshin Andreas
- 13:07 MST: The SRRL BMS Global 90-West PSP was installed at 11:36 MST today. -Afshin Andreas
- 13:58 MST: A SRRL BMS Research Channel History page has been created, that lists time frames and various instruments that have been connected to these research channels. -Afshin Andreas
- 12:47 MST: Fixed a bug with MIDC's SOLPOS Generator, that caused data for February 29 to not appear in the output if the date range spanned February 29 during a leap year and the start and end years were equal. -Afshin Andreas
- 12:35 MST: The signal cables (instrument to power supply) for the SRRL BMS Yankee instruments were strapped down using tie wraps to the BMS table today @ 8:30 MST. This seemed to fix the noisy signal problems with the instrument observed with high sample rate data (1 second), when there was movement or vibrations on the instrument deck. However, this caused the instruments to drift significantly out of calibration. They will be included in the next BORCAL. -Afshin Andreas
- 13:24 MST: The following SRRL BSRN instruments (which were calibrated in BORCAL 2004-01) were swapped out between 13:11 to 13:22 MST: Global PSP and Direct NIP. Calib factors were updated in the logger at 13:22 MST. -Afshin Andreas
- 12:49 MST: The following SRRL BMS instruments (which were calibrated in BORCAL 2004-01) were swapped out between 12:21 to 12:45 MST: Global 90 North PSP, Global 90 South PSP, and Global Normal. Calib factors were updated in the logger at 12:45 MST. -Afshin Andreas
- 10:19 MST: June 2004 Monthly Data Sets & Reports for SRRL BMS have been generated. -Afshin Andreas
- 09:51 MST: June 2004 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 07:58 MST: The SRRL BMS YES TSP-1 & TSP-700 were moved to separate power supplies on July 8, 2004 @ 12:30 MST, in an effort to improve the quality of the instruments. -Afshin Andreas
- 15:23 MST: The following SRRL BMS instruments (which were calibrated in BORCAL 2004-01) were swapped out between approx 14:24 to 15:13 MST: Diffuse PSP, Diffuse ShadowBand PSP, Global RG780 PSP, and Direct RG780 NIP. Calib factors were updated in the logger at approx 15:13 MST. -Afshin Andreas
- 15:55 MST: The SRRL BMS Display has been upgraded to 6 second updates (on site display). -Afshin Andreas
- 12:54 MST: The SRRL BMS Display was upgraded to 12 second updates yesterday (visible on the LabWindows Run-Time at SRRL). Web users will continue to see the usual 1 minute updates, however the data will be based on 12 second instantanous samples instead of 1 minute averages, therefore the irradiance plot will show a higher time resolution. -Afshin Andreas
- 07:02 MST: The ARM RCS site was shutdown on June 25. There are no plans to restart data acquisition. -Afshin Andreas
- 15:19 MST: The SRRL BMS TSP-700 & TSP-1 were calibrated in BORCAL 2004-01. New calibration factors were entered into the logger today @ 10:45 MST. -Afshin Andreas
- 09:03 MST: As of 8:45 MST today, the MIDC / SRRL computer system clocks are now maintained on Universal Time (UT) instead of Coordinated Universal Time (UTC), where UT = UTC + ΔUT1. ΔUT1 is derived from observation only and is reported in this bullitin, where ΔUT1 = UT1R-UTC. -Afshin Andreas
- 11:29 MST: The MIDC Data Acquisition System was upgraded from the old SOLPOS to the new Solar Position Algorithm (SPA) today @ 12:25 MST. This affects sites that include solar position information in their datastream, which are SRRL BMS, ARM RCS, and ECSU. -Afshin Andreas
- 13:11 MST: The SRRL BMS Display has been updated to use the new Solar Position Algorithm (SPA) for calculating the Zenith, Azimuth, and Airmass. -Afshin Andreas
- 13:09 MST: ΔT has been added to the SRRL BMS data stream. This value is derived from observation only, more details are available in the BMS Instrument History page. -Afshin Andreas
- 12:23 MST: The SRRL BMS Instrument History page has been updated for the 6 vertical instruments that were installed yesterday. -Afshin Andreas
- 11:48 MST: The SRRL BMS Global 90 North, East, & South instruments (both PSP and LI-200) were installed yesterday @ 14:30 MST. Calibration factors were applied to data at 00:00 MST today. -Afshin Andreas
- 12:49 MST: Fixed a bug when creating a user-defined calculation on a daily plot. If the user entered a value that was outside the "allowable" limits for a site (such as 6000 or greater) it would consider the result of the calculation as missing (even if the result was within the limits). This has been corrected to allow the user to enter any value. -Afshin Andreas
- 15:24 MST: The SRRL BMS instrument mounting diagram for the deck and two towers (PDF file) has been updated to include changes that occured with the BMS 9 upgrade. -Afshin Andreas
- 13:27 MST: Instructions for setting your PC time to SRRL's NTP Server using Windows 2000 built-in support is now available. -Afshin Andreas
- 15:02 MST: May 2004 Monthly Data Sets & Reports for SRRL BMS have been generated. -Afshin Andreas
- 14:18 MST: May 2004 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 10:55 MST: SRRL's Schott RSP data stream format has changed. The RSP & "RSP Diagnostic" sites have been merged into a single datastream. Also the hourly outputs that used to exist in the diagnostic site is now included in the one minute data. All historical data have been modified for this change. -Afshin Andreas
- 08:35 MST: The NWTC M2 dew point sensor was fixed yesterday at 13:30 MST. This also corrects the relative humidity problem. -Afshin Andreas
- 14:08 MST: The NWTC M2 Global PSP was replaced with instrument S/N 31555F3 between 9:06 to 9:20 MST today. The calibration factor was updated in the logger program today @ 14:06 MST. -Afshin Andreas
- 15:04 MST: The ECSU instrument history page has been updated to include the changes. -Afshin Andreas
- 14:16 MST: The ECSU data stream has been upgraded, all data since April 1, 2004 has been converted to the new format. New measurements include: CR10 battery voltage; New calculations include:
- Diffuse (corrected for shadowband)
- Broadband AOD
- Zenith, Azimuth, & Airmass
- Dew Point & Wind Chill Temperature
- 08:55 MST: SRRL SkyCam Daily Animations are now available. -Afshin Andreas
- 13:53 MST: The SRRL GPS Time Java Clock has been updated. Also instructions for setting your PC time using Windows XP built-in support is now available. -Afshin Andreas
- 12:52 MST: Two more Research Channels have been added to the SRRL BMS datastream. -Afshin Andreas
- 12:42 MST: April 2004 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 12:22 MST: April 2004 Monthly Data Sets & Reports for SRRL BMS have been generated. -Afshin Andreas
- 09:35 MST: Download output limits were placed on the MIDC Daily Data and Site Comparison Data of approx. 100 MB (uncompressed) for a site with 1-minute data. -Afshin Andreas
- 14:58 MST: The SRRL Global 40-South LI-1800 was calibrated and returned to service today @ 14:50 MST. -Afshin Andreas
- 08:00 MST: The SRRL Global 40-South LI-1800 Spectroradiometer was removed at 8:00 MST today for calibration. -Afshin Andreas
- 11:54 MST: A 10-ft section was added to the SPMD ROHN 25-G WT/Met Tower on April 28th. The wind speed and direction sensors were rewired and operational at ~1205 MST. The wind speed and direction sensors are now at the 40-ft level and the wind turbine is at the 43-ft level. -Afshin Andreas
- 09:54 MST: The SRRL BMS Sun Follower III was reset today @ ~ 8:45 MST. The tracker was down since April 21 due to tripped GFI. -Afshin Andreas
- 13:05 MST: The SRRL BMS Yankee TSP-700 and TSP-1 are working as of approx 9:30 MST. -Afshin Andreas
- 13:04 MST: The SRRL BMS Sci-Tec #2 is operational as of approx 08:50 MST. The AOCS is also running, however the Direct Quantum sensor does not seem to be working, problem is unknown at this time. -Afshin Andreas
- 14:42 MST: The SRRL BMS Yankee TSP-1 and TSP-700 are still down due to the power failure, problem is unknown at this time. -Afshin Andreas
- 14:13 MST: The SRRL AOCS has been shutdown due to Sci-Tec tracker problem. -Afshin Andreas
- 14:07 MST: The SRRL BMS Sci-Tec Tracker #2 is currently down due to signal cables getting hung up after power was restored to the tracker, affected instruments are: Direct MS-110, AOCS, Diffuse CM22, Diffuse UVB-1, Shaded CG4, Global LI-190. -Afshin Andreas
- 14:07 MST: There has been power outage at SRRL from 4:33 MST to 9:13 MST, due to a rat getting fried inside a transformer on the Mesa Top. -Afshin Andreas
- 11:46 MST: The SRRL BMS Sun Follower III tracker was reset on April 19 @ 7:00 MST. The tracker was down all day on the 18th. -Afshin Andreas
- 11:43 MST: There is still a problem with the NWTC M2 dew point sensor, which is causing the RH to read a constant 9-11%. The problem is being investigated. -Afshin Andreas
- 10:59 MST: The NWTC M2 MetOne dewpoint sensor DP200-B at 2m was replaced with a new unit at 10:45 MST April 15, 2004 -Pete Gotseff
- 07:21 MST: The SRRL AOCS now outputs calculated fields during all daylight hours. Prior to today, the software would only perform calculations when the air mass was less than 10. -Afshin Andreas
- 08:00 MST: There has been a problem with the NWTC M2 Dew Point Sensor (which is also used to calculate the Relative Humidity). The problem seems to have started on April 2. The RH does not read as high as it should. -Afshin Andreas
- 10:48 MST: The AOCS Instrument History page has been updated with new calibration factors and additional instrument information. -Afshin Andreas
- 09:42 MST: The Calibration factors for the (3) AOCS Quantum sensors were updated yesterday @ 15:15 MST. The Calibration factors for the (3) AOCS Irradiance sensors were updated today @ 9:32 MST. -Afshin Andreas
- 08:50 MST: Two fields have been added to the SRRL AOCS output: Diffuse Quantum (calculated) and AOCS Temperature (oC). Prior to April 14, diffuse quantum will read missing or zero and the temperature will read the raw count output from the temperature sensor. -Afshin Andreas
- 10:05 MST: The SRRL AOCS Direct Normal Quantum Sensor is reading high. This is causing the calculated Quantum Diffuse to go negative during clear sky conditions. The Quantum Diffuse/Quantum Global ratio is used to calculate the AOCS parameter, therefore the AOCS parameter is missing. Therefore there is no calculated spectra for the AOCS during clear skys. The direct quantum sensor needs to be calibrated. -Afshin Andreas
- 15:07 MST: Calibrated Global, Direct, & Diffuse instruments were installed at ECSU on April 5 @ 16:00 MST. The direct tracker is also functional. The instrument history page has been updated. -Afshin Andreas
- 12:49 MST: The RG780 dome on the SRRL BMS Global PSP was replaced on March 30, between 15:55 and 16:05 MST. -Afshin Andreas
- 13:01 MST: The SRRL AOCS is working as of today @ 12:50 MST. It was found that the Tattletale datalogger kept getting corrupt (due to possible intermitten rs-232 communications trouble). After all connections were dis/re-connected at the AOCS head, the unit began working again. -Afshin Andreas
- 12:21 MST: March 2004 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 12:02 MST: March 2004 Monthly Data Sets & Reports for SRRL BMS have been generated. -Afshin Andreas
- 14:23 MST: The SRRL AOCS has been down since April 3, due to the unit being plugged into the wrong GFI AC outlet (which tripped due to the SF3). The AOCS is still not functioning, problem is unknown at this time. -Afshin Andreas
- 08:52 MST: The RG780 Filter on the SRRL BMS NIP appears to have been replaced on March 31 at 10:40 MST. The new filter is now giving irradiance levels about 50% higher than the old filter. -Afshin Andreas
- 08:47 MST: ECSU is currently under maintenace. The Direct, Diffuse, & Global instruments were removed on 4/4/2004 @ 10:35 EST. -Afshin Andreas
- 08:42 MST: The SRRL BMS Sun Follower III was down since April 3, due to a tripped GFI. The tracker was reset today at 8:17 MST. This affected the following direct normal instruments: NIP #1, NIP RG780, CUVA2, CUVB2, LI-201. -Afshin Andreas
- 12:20 MST: The MIDC Web Server and/or data updates has been down or intermitten since March 22, due to a problem with an external Jaz 1 GB drive. The Jaz drive has been removed from the system and all data are up to date as of ~9:30 MST today. Data was only lost from the SRRL Global 40-South LI-1800 from 6:00 to 6:35 MST on March 24. -Afshin Andreas
- 16:14 MST: The SRRL BMS Global TUVR has been having signal drop-out problems again since March 12. -Afshin Andreas
- 10:11 MST: The SRRL BMS Direct TUVR is still having signal drop out problems. -Afshin Andreas
- 10:47 MST: The SRRL BMS Direct TUVR was shorted between 9:42 to 10:10 MST and the Global TUVR between 10:11 to 10:42 MST, in an effort to "reset" the instruments. -Afshin Andreas
- 09:40 MST: February 2004 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 09:28 MST: February 2004 Monthly Data Sets & Reports for SRRL BMS have been generated. -Afshin Andreas
- 09:10 MST: The SRRL BMS TUVRs (both Global and Direct) have been experiencing occasional signal drop outs over the past few days. -Afshin Andreas
- 14:07 MST: A new calibration factor was derived for the SRRL BMS Global TSP-1, based on 1-minute data taken February 14, 2004 from 9:26 to 17:00 MST, as compared with the calculated Global from the BMS Direct CH1 and Diffuse 8-48. The new calibration factor of 0.41233 W/m2/mV was entered into the logger today @ 13:57 MST. -Afshin Andreas
- 11:54 MST: Two "Research" channels are now available in the SRRL BMS data stream. These fields are for our research purposes and should be ignored by the user. -Afshin Andreas
- 08:49 MST: The SRRL BMS Diffuse Yankee UVB-1 was returned to service on Feb 10 @ 11:00 MST. The shading ball was not installed for the entire day, so that the instrument could be calibrated against the SRRL BMS Global UVB-1. The shade and new cal factor was installed at ~8:00 MST on Feb 11. -Afshin Andreas
- 15:01 MST: The SRRL BMS Monthly QA Field Definitions and QA Tests pages have been updated to reflect the BMS 9 upgrade. -Afshin Andreas
- 13:45 MST: January 2004 Monthly Data Sets & Reports for SRRL BMS have been generated. The following changes occured:
- Due to the BMS upgrade, 14 additional outputs (w/ flags) were added to the end of the monthly QA'd data file and well as 14 additional reports.
- The Flag Summary Table now has an additional Other Column.
- 12:53 MST: January 2004 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 08:06 MST: The BMS TWC 500nm Photometer still has a signal problem. -Afshin Andreas
- 13:54 MST: The BMS TWC 500nm Photometer was returned to service today @ ~14:43 MST. -Afshin Andreas
- 09:57 MST: The serial number for the SRRL BMS Global TUVR was not updated when the instrument was swapped in January 2003 with a S/N 18180. All BMS reports for 2003 have been corrected to reflect the correct S/N. -Afshin Andreas
- 14:28 MST: The BMS TSP-1 has been reading 25% high since Jan 9, problem is unknown at this time. -Afshin Andreas
- 13:49 MST: The BMS Live Display now shows the EKO MS110 500nm Turbidity instead of the TWC 500nm, due to the instrument problems. Note: The MS110 is also mounted on the more reliable Sci-Tec tracker. -Afshin Andreas
- 13:39 MST: There has been an intermitting problem with the BMS TWC 500nm Photometer since Jan 8, 2004. Instrument signal cable and power supply seem okay, problem is unknown at this time. -Afshin Andreas
- 14:31 MST: The SRRL BMS Instrument History page has been updated for changed that were made with BMS v9.0 (excluding instruments that have not been installed yet). -Afshin Andreas
- 10:21 MST: A new calibration factor was derived for the SRRL BMS Global TSP-700, based on 1-minute data taken December 27, 2003 from 10:45 to 12:00 MST, as compared with the calculated Global from the BMS Direct CH1 and Diffuse 8-48. The new calibration factor of 0.3823 W/m2/mV was entered into the logger today @ 8:48 MST. -Afshin Andreas
- 10:24 MST: The new SRRL BMS Global TSP-700 is reading about 100 W/m2 too low (using the factory calibration factor of 0.2825 W/m2/mV), problem is unknown at this time. -Afshin Andreas
- 09:12 MST: As part of the BMS 9.0 upgrade, an AM 416 MUX was added to the CR10X at the Radiometer Tower. All instruments at the Radiometer Tower (except the Upwelling PIR) are now connected to this AM 416. -Afshin Andreas
- 09:09 MST: The SRRL BMS Upwelling Shortwave PSP was returned to service on Jan 7 @ 14:44 MST. -Afshin Andreas
- 09:06 MST: SRRL BMS was upgraded to version 9.0 on January 6 @ 14:27 MST. All data from January 1, 2004 was converted to the new data stream. The following instruments were added:
- Global TSP-700 (installed 1/1 @ 0:00 MST)
- Upwelling Shortwave LI-200 (installed 1/7 @ 14:45 MST)
- Upwelling Quantum LI-190 (installed 1/7 @ 14:46 MST)
- Global 90-North PSP & LI-200 (not installed)
- Global 90-East PSP & LI-200 (not installed)
- Global 90-South PSP & LI-200 (not installed)
- Global 90-West PSP & LI-200 (not installed)
- Dew Point Temp (Tower)
- Albedo (LI-200)
- Albedo (LI-190)
- 16:33 MST: SRRL BMS has been upgraded to BMS 9.0, some instruments are not available yet, details to follow later. -Afshin Andreas
- 10:01 MST: December 2003 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 09:49 MST: December 2003 Monthly Data Sets & Reports for SRRL BMS have been generated. -Afshin Andreas
- 11:32 MST: The SRRL BMS Yankee TSP-1 started working again on Dec 24, 2003 at 12:10 MST, after another TSP-700 was connected to the same power supply. -Afshin Andreas
News Archives:
[1997] [1998] [1999] [2000] [2001] [2002] [2003] [2004] [2005] [2006] [2007] [2008] [2009] [2010] [2011] [2012] [2013] [2014] [2015] [2016] [2017] [2018] [2019] [2020] [2021] [2022] [2023]